# Optimizing Sparse SYK

### v2

@@ Acknowledgments
This work was primarily done while M.D. was visiting Caltech, and they would like to thank John Preskill and the Caltech SURF program for their support. E.R.A. was funded in part by the Walter Burke Institute for Theoretical Physics at Caltech.
We would also like to thank multiple anonymous reviewers for their helpful comments and suggestions.
@@ Appendix A Improved maximum eigenvalue upper bound
In this section we show a stronger upper bound on the maximum eigenvalue of sparse SYK Hamiltonians. [26] showed that the maximum eigenvalue of the SYK-4 model can be upper-bounded by $O(\sqrt{n})$ with probability exponentially close to 1 using a straightforward application of the matrix Chernoff bound, given as follows:
Proposition A.1 (Theorem 1.2 of [44]).
@@
We reproduce the proof of [26] for the dense SYK model for completeness. Since $\gamma_{I}$ in the Weyl–Brauer representation is a $2^{n}$ dimensional matrix and $\left\|\sum_{I}\gamma_{I}^{2}\right\|_{\mathrm{op}}=\left\|\sum_{I}\mathbb{I}\right\|_{\mathrm{op}}=\binom{2n}{4}$, we have:
$\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\binom{2n}{4}^{-1/2}\sum_{I}J_{I}\gamma_{I}\right)\geq C\sqrt{n}\right)$ $\leq\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\sum_{I}J_{I}\gamma_{I}\right)\geq Cn^{5/2}\right)$
@@
$\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\binom{2n}{4}^{-1/2}p^{-1/2}\sum_{I}X_{I}J_{I}\gamma_{I}\right)\geq C\sqrt{n}\right)$ $\leq\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\sum_{I\subseteq\mathrm{supp}(\vec{x})}J_{I}\gamma_{I}\right)\geq Cn^{5/2}p^{1/2}\right)$
@@
For large enough constant $C$ the right hand side scales as $\exp(-\Omega(n))$. Using a union bound, we can upper-bound the probability of the event that, (1) either the number of non-zero entries in our binomial vector is greater than twice the expected value, or (2) that the maximum eigenvalue is greater than $C\sqrt{n}$ given the number of non-zero entries in our binomial vector is less than twice the expected value, as less than $\exp(-\Omega(n))$. Formally,
$\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}(H_{SSYK})\geq C\sqrt{n}\right)$ $\leq\operatorname*{\mathbb{P}}\left(\left(\lambda_{\mathrm{max}}(H_{SSYK})\geq C\sqrt{n}\;\Big|\;|\vec{x}|_{0}<2\operatorname*{\mathbb{E}}[|\vec{x}|_{0}]\right)\vee\left(|\vec{x}|_{0}\geq 2\operatorname*{\mathbb{E}}[|\vec{x}|_{0}\right)\right)$
@@ Appendix B Energy upper bounds on low-magic fermionic ansatz
The tensor network construction of [25] was originally devised to prove upper bounds on the energy achievable by two different ansatz for the SYK model. However the proofs apply to any Hamiltonians which satisfy the norm bound of $O(k)^{2k}$ given in Lemma˜4.2 when $p\geq\Omega(\log n/n^{2})$.
Superpositions of Gaussian states
@@ Appendix C Moment bound proofs
Proposition C.1 (Decoupling of random polynomials [18]).
$\operatorname*{\mathbb{P}}\left[\left|\sum_{1\leq i_{1}\neq\ldots\neq i_{k}\leq n}f_{i_{1}\ldots i_{k}}(Y_{i_{1}},\ldots,Y_{i_{k}})\right|\geq t\right]\leq C_{k}\operatorname*{\mathbb{P}}\left[\left|\sum_{1\leq i_{1}\neq\ldots\neq i_{k}\leq n}f_{i_{1}\ldots i_{k}}(Y_{i_{1}}^{(1)},\ldots,Y_{i_{k}}^{(k)})\right|\geq t\right],$ (81)
@@
Previous works on optimizing SYK [26, 28] used this decoupling property when the $Y_{i}$ were Gaussian random variables. However, the property more generally applies when the $X_{i}$ are drawn independently from any distribution [18].
@@ C.1 Bounding $D_{0}$
where the $(\cdot)^{\left(i\right)}$ superscripts denote independent (“decoupled”) random variables, the $X$ variables are independent $\text{Bernoulli}\left(p\right)$ random variables, and the summation uses the index notation from [26]: $r,s,t,u,w\in\left[n_{1}\right]$ and $a,b,d\in\left[n_{2}\right]$. In the above, the superscripts indicate that the random variables are independent if the superscript index is not equal. This expression is obtained by expanding the $S=(S_{1},S_{2},S_{3})$ sets into individual indices, because $f(S,S^{\prime},S^{\prime\prime},j,k,l)=1$ if $(|S^{\prime\prime}\cap S|\text{ is odd})\wedge(|S^{\prime}\cap(S^{\prime\prime}\triangle S)|+\delta_{k,l}\text{ is odd})$ forces at least two pairs of the 12 indices to be identical in the non-zero terms of the summation. Thus, due to symmetry, we can arbitrarily fix two pairs of indices to be equivalent ($u$ and $t$ respectively) and sum over all remaining indices, and the set of terms in $D_{0}$ in our actual expression will be a subset of the summation in Eq.˜82. This will only increase the $2n^{th}$ moment.
@@
Let $B_{\text{small}}$ denote the event $\left\{\forall t,u\ B^{(1)}_{t}(C_{1}n^{3},p)\leq(C_{1}+c)pn^{3}\wedge B^{(3)}_{u}(C_{3}n^{3},p)\leq(C_{3}+c)pn^{3}\right\}$ for some sufficiently large constant $c$, and denote $B_{\text{large}}$ as its complement. We have:
@@
We can directly use [26] to bound the first term. As a reminder, we set $\tilde{D}_{0,\text{norm}}\triangleq\tilde{D}_{0}/\frac{Cp^{-3/2}}{n^{5}}$. We have:
@@
where $\left\|\sum_{tu}H_{t}K_{tu}L_{u}\right\|_{2n}^{2n}\leq O(n^{3/2})$ is a bound due to [26]. By the union bound,
@@
Let $C_{i}^{+}\triangleq C_{i}+c$ for $i=1,2,3$. By a union bound and Chernoff bound on the binomial random variables (Eq.˜7),
$\operatorname*{\mathbb{P}}(B_{\text{large}})$ $=\operatorname*{\mathbb{P}}\left(\bigcup_{t,u}\left(B_{t}(C_{1}n^{3},p)>C_{1}^{+}pn^{3}\right)\lor\left(B_{tu}(C_{2}n^{2})>C_{2}^{+}pn^{2}\right)\lor\left(B_{u}(C_{3}n^{3},p)>C_{3}^{+}pn^{3}\right)\right)$
@@ C.2 Bounding $D_{1}$
To bound this, we can use hypercontractivity (Bonami’s lemma [9]). $D_{1}$ is a degree-3 Gaussian polynomial with $B=\text{Binomial}\left(O(n^{4}),p\right)$ terms. Let $B_{\text{small}}$ denote the event $B\leq(1+c)pn^{4}$ for some constant $c$, and $B_{\text{large}}$ as its complement. We have:
@@ C.3 Bounding $D_{2},D_{3},D_{4}$
to show that $\operatorname*{\mathbb{E}}[\tilde{D}_{2,\text{norm}}^{2n}\mid B_{\text{small}}]\leq C^{2n}n^{11n}p^{3n}$. The Laurent–Massart [37] lower-bound for chi-squared random variables gives:
@@
Finally to show the correctness of Eq.˜99,
@@
Assume every binomial variable $B$ takes its maximum value. From the dense case [26], we have:

### v3

@@ Acknowledgments
This work was primarily done while M.D. was visiting Caltech through the Summer Undergraduate Research Fellowship (SURF), funded in part by the Institute for Quantum Information and Matter at Caltech, an NSF Physics Frontiers Center (PHY-2317110). They would like to thank John Preskill and the Caltech SURF program for their continued and invaluable support.
M.D. and R.K. acknowledge funding from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Center, Quantum Systems Accelerator (DE-SCL0000121). E.R.A. was funded in part by the Walter Burke Institute for Theoretical Physics at Caltech.
The authors would also like to thank multiple anonymous reviewers for their helpful comments and suggestions.
@@ Appendix A Improved maximum eigenvalue upper bound
In this section we show a stronger upper bound on the maximum eigenvalue of sparse SYK Hamiltonians. [undeff] showed that the maximum eigenvalue of the SYK-4 model can be upper-bounded by $O(\sqrt{n})$ with probability exponentially close to 1 using a straightforward application of the matrix Chernoff bound, given as follows:
Proposition A.1 (Theorem 1.2 of [undefaq]).
@@
We reproduce the proof of [undeff] for the dense SYK model for completeness. Since $\gamma_{I}$ in the Weyl–Brauer representation is a $2^{n}$ dimensional matrix and $\left\|\sum_{I}\gamma_{I}^{2}\right\|_{\mathrm{op}}=\left\|\sum_{I}\mathbb{I}\right\|_{\mathrm{op}}=\binom{2n}{4}$, we have:
$\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}(H_{SYK})\geq C\sqrt{n}\right)$ $\leq\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\sum_{I}J_{I}\gamma_{I}\right)\geq Cn^{5/2}\right)$
@@
$\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}(H_{SSYK})\geq C\sqrt{n}\right)$ $\leq\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}\left(\sum_{I\subseteq\mathrm{supp}(\vec{x})}J_{I}\gamma_{I}\right)\geq Cn^{5/2}p^{1/2}\right)$
@@
For large enough constant $C$ the right hand side scales as $\exp(-\Omega(n))$. Using a union bound, we can upper-bound the probability of the event that, (1) either the number of non-zero entries in our binomial vector is greater than twice the expected value, or (2) that the maximum eigenvalue is greater than $C\sqrt{n}$ given the number of non-zero entries in our binomial vector is less than twice the expected value, as less than $\exp(-\Omega(n))$. Formally, $\operatorname*{\mathbb{P}}\left(\lambda_{\mathrm{max}}(H_{SSYK})\geq C\sqrt{n}\right)$ is upper-bounded by
$\operatorname*{\mathbb{P}}\left(\left(\lambda_{\mathrm{max}}(H_{SSYK})\geq C\sqrt{n}\;\Big|\;|\vec{x}|_{0}<2\operatorname*{\mathbb{E}}[|\vec{x}|_{0}]\right)\vee\left(|\vec{x}|_{0}\geq 2\operatorname*{\mathbb{E}}[|\vec{x}|_{0}\right)\right)$
@@ Appendix B Energy upper bounds on low-magic fermionic ansatz
The tensor network construction of [undefw] was originally devised to prove upper bounds on the energy achievable by two different ansatz for the SYK model. However the proofs apply to any Hamiltonians which satisfy the norm bound of $O(k)^{2k}$ given in Lemma˜4.2 when $p\geq\Omega(\log n/n^{2})$.
Superpositions of Gaussian states.
@@ Appendix C Moment bound proofs
Proposition C.1 (Decoupling of random polynomials [undefav]).
$\operatorname*{\mathbb{P}}\left[\left|\sum_{1\leq i_{1}\neq\ldots\neq i_{k}\leq n}f_{i_{1}\ldots i_{k}}(Y_{i_{1}},\ldots,Y_{i_{k}})\right|\geq t\right]\leq C_{k}\operatorname*{\mathbb{P}}\left[\left|\sum_{1\leq i_{1}\neq\ldots\neq i_{k}\leq n}f_{i_{1}\ldots i_{k}}(Y_{i_{1}}^{(1)},\ldots,Y_{i_{k}}^{(k)})\right|\geq t\right]$ (80)
@@
Previous works on optimizing SYK [undeff, undefs] used this decoupling property when the $Y_{i}$ were Gaussian random variables. However, the property more generally applies when the $X_{i}$ are drawn independently from any distribution [undefav].
@@ C.1 Bounding $D_{0}$
where the $(\cdot)^{\left(i\right)}$ superscripts denote independent (“decoupled”) random variables, the $X$ variables are independent $\text{Bernoulli}\left(p\right)$ random variables, and the summation uses the index notation from [undeff]: $r,s,t,u,w\in\left[n_{1}\right]$ and $a,b,d\in\left[n_{2}\right]$. In the above, the superscripts indicate that the random variables are independent if the superscript index is not equal. This expression is obtained by expanding the $S=(S_{1},S_{2},S_{3})$ sets into individual indices, because $f(S,S^{\prime},S^{\prime\prime},j,k,l)=1$ if $(|S^{\prime\prime}\cap S|\text{ is odd})\wedge(|S^{\prime}\cap(S^{\prime\prime}\triangle S)|+\delta_{k,l}\text{ is odd})$ forces at least two pairs of the 12 indices to be identical in the non-zero terms of the summation. Thus, due to symmetry, we can arbitrarily fix two pairs of indices to be equivalent ($u$ and $t$ respectively) and sum over all remaining indices, and the set of terms in $D_{0}$ in our actual expression will be a subset of the summation in Eq.˜81. This will only increase the $2n^{th}$ moment.
@@
Let $B_{\text{small}}$ denote the event: $\forall t,u\ B^{(1)}_{t}(C_{1}n^{3},p)\leq(C_{1}+c)pn^{3}\wedge B^{(3)}_{u}(C_{3}n^{3},p)\leq(C_{3}+c)pn^{3}$ for some sufficiently large constant $c$, and denote $B_{\text{large}}$ as its complement. We have:
@@
We can directly use [undeff] to bound the first term. As a reminder, we set $\tilde{D}_{0,\text{norm}}\triangleq\tilde{D}_{0}/\frac{Cp^{-3/2}}{n^{5}}$. We have:
@@
where $\left\|\sum_{tu}H_{t}K_{tu}L_{u}\right\|_{2n}^{2n}\leq O(n^{3/2})$ is a bound due to [undeff]. By the union bound,
@@
Let $C_{i}^{+}\triangleq C_{i}+c$ for $i=1,2,3$. By a union bound and Chernoff bound on the binomial random variables (Eq.˜7), $\operatorname*{\mathbb{P}}(B_{\text{large}})$ is bounded by
$\operatorname*{\mathbb{P}}\left(\bigcup_{t,u}\left(B_{t}(C_{1}n^{3},p)>C_{1}^{+}pn^{3}\right)\lor\left(B_{tu}(C_{2}n^{2})>C_{2}^{+}pn^{2}\right)\lor\left(B_{u}(C_{3}n^{3},p)>C_{3}^{+}pn^{3}\right)\right)$
@@ C.2 Bounding $D_{1}$
To bound this, we can use hypercontractivity (Bonami’s lemma [undefaw]). $D_{1}$ is a degree-3 Gaussian polynomial with $B=\text{Binomial}\left(O(n^{4}),p\right)$ terms. Let $B_{\text{small}}$ denote the event $B\leq(1+c)pn^{4}$ for some constant $c$, and $B_{\text{large}}$ as its complement. We have:
@@ C.3 Bounding $D_{2},D_{3},D_{4}$
to show that $\operatorname*{\mathbb{E}}[\tilde{D}_{2,\text{norm}}^{2n}\mid B_{\text{small}}]\leq C^{2n}n^{11n}p^{3n}$. The Laurent–Massart [undefap] lower-bound for chi-squared random variables gives:
@@
Finally to show the correctness of Eq.˜98,
@@
Assume every binomial variable $B$ takes its maximum value. From the dense case [undeff], we have: